OpenStudy (loser66):

\(f(x) =\begin{cases}h(x) ~~~if~~~ x \in \mathbb Q\\g(x) ~~~if~~~x\notin \mathbb Q\end{cases}\) Prove that f(x) is differentiable at x = a if and only if g(a) =h(a) and g'(a) =h'(a) Please, help

OpenStudy (loser66):

@xapproachesinfinity

OpenStudy (loser66):

My attempt: suppose f(x) is diff. at x =a, then f'(x) is defined and \(f'(a) = lim_{x\rightarrow a}\dfrac{f(x)-f(a)}{x-a}\)

OpenStudy (loser66):

since f(x) is a piecewise function, \(f'(x) =\begin{cases}h'(x) ~~~if~~x\in\mathbb Q\\g'(x) ~~~if ~~~x\notin \mathbb Q\end{cases}\)

OpenStudy (loser66):

Moreover, f(x) is diff. at x=a, hence f(x) is continuous at x =a, that is h(x) = h(a)

OpenStudy (xapproachesinfinity):

yeah but that what we are trying to prove here if h'(a)=g'(a) then f must be diff at x=a

OpenStudy (loser66):

and g(x) = g(a) AND h(a) =g(a)

OpenStudy (xapproachesinfinity):

yes the first is the continuity case well f diff at means that g(a)=h(a)

OpenStudy (loser66):

@xapproachesinfinity we go both ways, right? I am proving --> direction

OpenStudy (xapproachesinfinity):

iff and only mean both ways

OpenStudy (xapproachesinfinity):

so are you trying contradiction?

OpenStudy (loser66):

Nope, why? I go directly.

OpenStudy (xapproachesinfinity):

hmm you made a supposition so i thought you want to go with contradiction

OpenStudy (loser66):

f is diff. hence, f is conti. at x =a , that is h(x) = g(x) at a. It means h(a) =g(a)

OpenStudy (xapproachesinfinity):

i'm not sure that is a good proof hehe

OpenStudy (xapproachesinfinity):

what if we start backwards?

OpenStudy (loser66):

Let finish this way, please

OpenStudy (loser66):

we just have h(a) = g(a) , still need h'(a) =g'(a)

OpenStudy (xapproachesinfinity):

well if f is diff on a and since it is peice-wise then h and g must be diff as well and the only that can happen is when h'(a)=g'(a) is there are not equal then our assumption must be wrong

OpenStudy (xapproachesinfinity):

if there are not*

OpenStudy (xapproachesinfinity):

that is way i think the proof is weak lol

OpenStudy (loser66):

ok, kid, I have a proposition for that. It says: f( x) is diff. at x =a iff \(h'|_Q(a) = g'|_{R\Q}(a)\)

OpenStudy (loser66):

hence h'(a) =g'(a)

OpenStudy (loser66):

Now, other way.

OpenStudy (xapproachesinfinity):

hmm is that R\Q ?

OpenStudy (loser66):

\(\notin \mathbb Q= \mathbb R/\mathbb Q\)

OpenStudy (xapproachesinfinity):

yep i know i'm just making sure that's what you mean i didn't know such proposition if that's true then your proof makes some sense

OpenStudy (xapproachesinfinity):

the other way we prove that both h and g are differential at a an equal and that there a continuity with both conditions we can say f is diffrentiable

OpenStudy (loser66):

By definition \(h'(a) = lim_{x\rightarrow a}\dfrac{h(x) -h(a) }{x-a} =lim_{x\rightarrow a}\dfrac{g(x) -g(a) }{x-a}= g'(a)\)

OpenStudy (xapproachesinfinity):

that also means there both continuous at x=a so lim h(x)=lim (g(x))

OpenStudy (xapproachesinfinity):

they are*

OpenStudy (loser66):

and.... oohhooh... old man is stuck.

OpenStudy (xapproachesinfinity):

well if you have h and g diff at a then that implies they are also continuous at a so g(a)=h(a)

OpenStudy (xapproachesinfinity):

to be honest i feel there is messing links haha

OpenStudy (loser66):

hey kid, I saw @kmeis002 is typing but why it takes toooo long to pop up?? hehehe

OpenStudy (xapproachesinfinity):

i was gonna ask the same thing lol perhaps he/she has something to say :)

OpenStudy (loser66):

hehehe, let me tag someone else but it is soooooo impolite when other is helping me.

OpenStudy (xapproachesinfinity):

hehehe why impolite? tag jim

OpenStudy (kmeis002):

Let \(x = a\) where \( a \in \mathbb{Q} \) then \( f(a) = h(a) \). For \( f\) to be continous at a, for any open ball \(f(B_\delta(a)) \subset B_\epsilon (f(a)) \). Assume \(g(a) \neq h(a)\). Since the reals are dense then \(\exists b \in B_\delta(a) \quad \text{s.t.} \quad b \notin \mathbb{Q}\). This implies we can choose an open ball of some radius such that \(\exists b \: \text{s.t.} \:f(b) \notin B_\epsilon(f(a)) \quad \text{and} \quad b \in B_\delta(a) \). Therefore, its not continuous. I believe you can show the same for the case of the derivatives to show the derivatives are continuous at a only if they are equal. Sorry, just trying to formulate, and multi-task. So it may not be correct at the moment, just throwing it out there.

OpenStudy (loser66):

ohoh... he is playing with balls.

OpenStudy (xapproachesinfinity):

what are the balls? no idea here lol

OpenStudy (loser66):

kid, you don't know this stuff, it is from advance cal

OpenStudy (xapproachesinfinity):

darn i feel bad now :)

OpenStudy (kmeis002):

Open balls are collection of numbers based on a radius r. Just an open set around a point.

OpenStudy (xapproachesinfinity):

hmm kinda heard something similar in algebra or something

OpenStudy (loser66):

@kmeis002 question, if \(b\in B_\delta (a)\) and \(a\in \mathbb Q\), then b can be in Q , can be not in Q also. How can you assume that b is not in Q?

OpenStudy (kmeis002):

For any open ball around a rational, there exists some element that is not rational.

OpenStudy (loser66):

because R is so dense

OpenStudy (loser66):

oh, got you \(\exists\), not all

OpenStudy (loser66):

but it goes nowhere, right?

OpenStudy (kmeis002):

Nowhere, how so? I'm a bit confused at this question

OpenStudy (kmeis002):

I admit, my post isn't fully complete, I'm pondering at the moment.

OpenStudy (loser66):

we need prove if f is diff. then f is conti. to get h(a) = g(a) . YOur balls give me f is NOT conti. hence.... I... stuck.

OpenStudy (kmeis002):

Mainly the explanation of the fact that at \(b, f(b) = g(b)\). Then we just need to show that we arrive at the contradiction (not yet complete). My post is hinting that if we assume the opposite, the function is not continuous. The same argument (incomplete) can be constructed for \(f'(a)\), which may take care of the differentiable. In fact, that would be all you would need to show since that implies continuity at a.

OpenStudy (loser66):

Ok, I am waiting, please, complete the stuff.

OpenStudy (kmeis002):

I'll see what I can do, if its anything at all. I apologize if my replies are not in the most timely matter. I have to take care of something for a moment, but if rigorous explanation arises I will share it.

OpenStudy (loser66):

Thank you for the help. :)

OpenStudy (kmeis002):

Alright, bear with me, but I think I have a half decent explanation to follow. If its not that great, I apologize. My formal training is not in pure mathematics so my proof skills go as far as my hobby does.

OpenStudy (kmeis002):

(Im am going to post these piecemeal so you dont have to wait for one long text) First we have to assume that \(g'(x), h'(x)\) are continuous about \(a\). And as discussed, if we can show f'(a) is continuous, then f(a) is continuous. A similar argument can be made for the \(g(a) = h(a)\) condition, if you please. Assume the opposite is true \(g'(a) \neq h'(a)\). This means \(|g'(a) -h'(a)| = \xi >0 \). Since \(g'(a), h'(a)\) are continuous about \(a\), there must exist an interval of radius \(\delta\), \( \left ( B_\delta(a) \right ) \) such that any value in the interval \( \alpha \in B_\delta(a) \) has the property \(|g'(\alpha) -h'(\alpha)| > 0\). Basically, we can choose an interval where \(g'(x) \neq h'(x)\) for some interval by the definition of continuity. Since the reals are dense, there must exist a \(b \in B_\delta(a) \) such that \(b \in \mathbb{R}\setminus \mathbb{Q}\)

OpenStudy (kmeis002):

If we assume \(a \in \mathbb{Q}\) then at \(a\) we get that \(f'(a) = h'(a)\) and at the \(b\) we know that \(f'(b) = g'(b) \). In fact, we know that since \(|g'(a)-h'(a)|>0\) we can also that that the arbitrary interval can be chosen to have the property \(|g'(b) - h'(a)| > 0 \) (essentially from the Intermediate Value Theorem, if g(a) and h(a) are not equal, one is greater and one is less than on some interval until they equal each other. That is if they ever do). This property will help us show that \(f'(a)\) is not continuous and therefore \(f(a)\) is not differentiable at a.

OpenStudy (kmeis002):

Finally, from the epsilon delta definition of a limit and since \(f'(b) = g'(b)\) and \( f'(a) = h'(a)\), we can show that \(|g'(b) -b'(a)| =|f'(b) -f'(a)| >0 \). Since \(b\) is some arbitrary irrational number in the interval, it can be expressed as \(b\) as any \( x \in B_\delta(a) \) that is irrational. Then \(|f'(x) -f'(a)| >\epsilon > 0\). That is, we can choose a \(b\) to make the difference around \(f'(a)\) not arbitrarily small. Infact, the smaller the interval, the closer it is to \(\xi\) (a constant value defined by \(g'(a), h'(a)\). This shows that \(f'(x)\) is not continuous about \(a\) if \(g'(a) \neq h'(a) \). Proving your proposition you posted above. I think with some effort, it can be formalized and made more compact, but that is where my thoughts were going. I wanted to try to write them in a more intuitive way.

OpenStudy (loser66):

Question: how |g'(a) -h'(a)|>0?? is it not that we assume that g'(a) = h'(a)?

OpenStudy (kmeis002):

The proof is by contradiction, in a sense. We assume \(g'(a) \neq h'(a)\) and show that this produces a discontinuous \(f'(a)\), implying that \(f'(a)\) is continuous iff \(g'(a) = h'(a)\)

OpenStudy (loser66):

oh, yeah,

OpenStudy (kmeis002):

My explanation got hairy on the third post. I got sloppy with the limit and differed to an English explanation. I think you can show that for some \(0<|x-a|<\delta\) produces a \(|f'(x)-f(a)|=\xi>\epsilon \)

OpenStudy (kmeis002):

*deferred

OpenStudy (kmeis002):

Thank you for the exercise!

OpenStudy (loser66):

Much appreciate :)